# InstructRAG: Instructing Retrieval-Augmented Generation with Explicit Denoising

### v1

@@ Appendix A Implementation Details
Training details. Our models are trained on 4 Nvidia H100 GPUs with 80GB memory via full-parameter fine-tuning. We use fully sharded data parallelism (FSDP) for distributed training, along with FlashAttention [11] and bf16 mixed precision training enabled for computation efficiency. By default, all models are trained using the Adam optimizer [38] for 2 epochs, with a batch size of 128, a learning rate of 2.5e-5, and a cosine learning rate schedule with 3% warmup steps. For the trainable baseline vanilla SFT, we use a slightly different learning rate of 2e-5 based on our hyper-parameter search results. To fairly compare with Self-RAG, we re-implement it using Llama-3-Instruct-8B. We also optimize its performance through an extensive hyper-parameter search with learning rates in [8e-6, 1e-5, 2e-5] and training epochs in [1, 2, 3], and find that a learning rate of 1e-5 with a single training epoch leads to the best result. The maximum token length for all models is fixed at 4096.
@@
Appendix B Prompt Templates
@@
Inference prompts. Below we present the inference prompts for both training-free and trainable RAG methods used in this work, including in-context RALM (Table 10), few-shot CoT with demonstrations (Table 11), and vanilla supervised fine-tuning (Table 12). Note that for a fair comparison, the inference prompt for our InstructRAG-FT (Table 3) is exactly the same as vanilla SFT (Table 12). Similarly, the inference prompt for InstructRAG-ICL (Table 2) shares the same inference instruction as the few-shot CoT with demonstrations (Table 11). The only difference between the prompts of these two methods lies in the demonstrations where InstructRAG-ICL employs denoising question-rationale pairs, while few-shot CoT with demonstrations uses plain question-answer pairs.

### v2

@@ Appendix A Implementation Details
Training details. Our models are trained on 4 Nvidia H100 GPUs with 80GB memory via full-parameter fine-tuning. We use fully sharded data parallelism (FSDP) for distributed training, along with FlashAttention [14] and bf16 mixed precision training enabled for computation efficiency. By default, all models are trained using the Adam optimizer [45] for 2 epochs, with a batch size of 128, a learning rate of 2.5e-5, and a cosine learning rate schedule with 3% warmup steps. For the trainable baseline vanilla SFT, we use a slightly different learning rate of 2e-5 based on our hyper-parameter search results. To fairly compare with Self-RAG and RetRobust, we re-implement them using Llama-3-Instruct-8B. We also optimize their performance through an extensive hyper-parameter search with learning rates in [8e-6, 1e-5, 2e-5] and training epochs in [1, 2, 3]. For Self-RAG, we use a learning rate of 1e-5 with a single training epoch. For RetRobust, we use a learning rate of 2e-5 with two training epochs. The only exception is the training for RetRobust on 2WikiMultiHopQA, where we train the model for 5 epochs on the augmented training set released by the original authors. The maximum token length for all models is fixed at 4096.
@@
Appendix B Case Study
Appendix C Prompt Templates
@@
Inference prompts. Below we present the inference prompts for both training-free and trainable RAG methods used in this work, including in-context RALM (Table 9), few-shot demonstrations with instruction (Table 10), and vanilla supervised fine-tuning (Table 11). Note that for a fair comparison, the inference prompt for our InstructRAG-FT is exactly the same as vanilla SFT. Similarly, the inference prompt for InstructRAG-ICL shares the same inference instruction as the few-shot demonstrations with instruction. The only difference between the prompts of these two methods lies in the demonstrations where InstructRAG-ICL employs denoising question-rationale pairs, while few-shot demonstrations with instruction uses plain question-answer pairs.